Exercise 10-5 Working Backwards from Labor Variances [LO10-2]

The auto repair shop of Quality Motor Company uses standards to control the labor time and labor cost in the shop. The standard labor cost for a motor tune-up is given below:

  Job Standard Hours   Standard Rate   Standard Cost  
  Motor tune-up 2.5         $9.00         $22.50       

The record showing the time spent in the shop last week on motor tune-ups has been misplaced. However, the shop supervisor recalls that 50 tune-ups were completed during the week, and the controller recalls the following variance data relating to tune-ups:

  Labor rate variance $ 87 F
  Labor spending variance $ 93 U
Required:
1. Determine the number of actual labor-hours spent on tune-ups during the week.
Answer #1

Labor efficiency variance

= 93 U - 87 F

= 180 U

(SH - AH) * SR = - 180

(2.5*50 - AH) * 9 = - 180

(125 - AH) =-180/9 = - 20

AH = 125 + 20

Actual hours = 145